OpenStudy (anonymous):

If a wave has amplitude of 2 meters, a wavelength of 2 meters, and a frequency of 10 Hz, and a period of 1 second, then at what speed is the wave moving? a. 40 m/s b. 20 m/s c. 10 m/s d. 5 m/s

OpenStudy (anonymous):

I think its b but i want to double check

OpenStudy (ladiesman217):

i got b as well good job

OpenStudy (anonymous):

Thanks

OpenStudy (ladiesman217):

no problem any other questions

OpenStudy (anonymous):

What is the period of vibration for a wave with a frequency of 150 hertz? a. 0.006 seconds b. 0.06 seconds c. 15 seconds d. 150 seconds

OpenStudy (anonymous):

dont need a direct answer just help solving it

OpenStudy (ladiesman217):

We use this equation: Period (T) = 1/Frequency (f) Basically, we substitute the values in to solve for period (T): T = 1/f T = 1/(150) then what does T equal after you divide it by f